OpenStudy (helpmeplzz):

what is the nth term of the sequence 1 2 4 8 16 32?

hartnn (hartnn):

do u realise that those are powers of 2 ? so the n'th term is 2^(n-1) ok?

OpenStudy (helpmeplzz):

okk thxxx

OpenStudy (helpmeplzz):

Anf how can i prove that

OpenStudy (helpmeplzz):

and

hartnn (hartnn):

ok, put n=1,2,3.....and see whether u get that sequence.......

OpenStudy (campbell_st):

ok... look at the common ratio by comparing the ratio of terms \[\frac{T_{2}}{T_{1}} = \frac{T_{3}}{T_{2}} .... \]

OpenStudy (campbell_st):

this will show that the series is geometric... A term in a geometric series is \[T_{n} = a r^{n -1}\] a = 1st term, r = is the common ratio just substitute

OpenStudy (helpmeplzz):

@campbell_st can you show me for the first term

OpenStudy (campbell_st):

1st term in the sequence is 1 the sequence is 1, 2, 4, 8, 16, ...

OpenStudy (helpmeplzz):

yes

OpenStudy (campbell_st):

the common ratio is 2... 2/1 = 4/2 = 8/4 = 16/8 ..... so the series is geometric hope it helps

OpenStudy (helpmeplzz):

okkk Thanks alot